OpenStudy (anonymous):

Given that measurement z is congruent to measurement x, measurement w = 8x+7, and measurement y = 5x+24, find the value of x for which wxyz must be a parallelogram.

OpenStudy (anonymous):



jhonyy9 (jhonyy9):

z=x w=8x+7 y=5x+24 - so if wxyz must be parallelogramm and we know that z=x so then there must to be w=y - if w=y than 8x+7=5x+24 --- 3x=17 --- x=17/3